OpenStudy (anonymous):

What about this one @Diyadiya Given the following perfect square trinomial, fill in the missing term. x^2 − 16x + ____

OpenStudy (anonymous):

1

OpenStudy (anonymous):

So it would be 8?

OpenStudy (diyadiya):

Do using the discriminant method b^2 - 4ac =0 (since it is a perfect square trinomial ) and find c

OpenStudy (diyadiya):

(-16)^2 -4*1*c =0

OpenStudy (diyadiya):

can you continue ?

OpenStudy (diyadiya):

What is (-16)^2 ?

OpenStudy (savvy):

clearly it would be something like (ax + b)^2 expand to get (a^2)(x^2) + b^2 + 2abx so a^2 = 1 and 2ab = 16 i.e. ab = 8 a= 1 or -1 so if a =1 then b=8 and if a = -1, b = -8 so it is (-x - 8)^2 or (x + 8)^2 also we can write the first as [(-1)^2](x+8)^2 which is (x+8)^2 so both the answers are same and the final answer is (x+8)^2

OpenStudy (anonymous):

256

OpenStudy (diyadiya):

256 - 4*1*c =0 256 - 4c =0 add 4c on both sides 256 - 4c +4c =4c 256 = 4c Now divide both sides by 4 what do you get?

OpenStudy (anonymous):

64

OpenStudy (diyadiya):

Yes 64 is the missing term :)

OpenStudy (anonymous):

Ohh.....

OpenStudy (anonymous):

So you square then divide.

OpenStudy (diyadiya):

Square ? No Use this formula b^2 - 4ac =0 (Since its PST ) Substitute the values of b , a & c (here we have to find c ,so let c remain like that )

OpenStudy (anonymous):

I see

OpenStudy (savvy):

@careless850 didn't you get my way...??

OpenStudy (diyadiya):

:)

OpenStudy (anonymous):

@Savvy No.

OpenStudy (anonymous):

Tooo complicated. i like the abc method.

OpenStudy (savvy):

ok.....do as u are comfortable with...

OpenStudy (anonymous):

:)